OpenStudy (anonymous):

l=8 w=4 h=2 Find the diagonal(d) of the base

OpenStudy (anonymous):

I believe this is a cuboid so base = l*b so base is a rectangle with length = 8 and width = 4 so it looks like  so Using Pythagoras Theorem Diagonal^2 = 8^2 + 4^2 Diagonal ^2 = 64 + 16 Diagonal^2 = 80 so Diagonal = root(80) = 4root(5)